Two balls are thrown simultaneously, A vertically upward with a speed of 20m/s from the ground , and B vertically downward from a height of 40m with the same speed along the same line of motion . At what point do two balls meet ?

1
Expert's answer
2021-12-02T10:02:57-0500

"y_1=0+v_0t-gt^2\/2"


"y_2=40-v_0t-gt^2\/2"


"y_1=y_2" . So, we have


"v_0t-gt^2/2=40-v_0t-gt^2/2\to t=40/(2v_0)=40/(2\cdot20)=1\ (s)"


"y=y_1=y_2=v_0t-gt^2/2=20\cdot1-9.8\cdot1^2/2=15.1\ (m)" . Answer